# Forecasting Clinical Risk from Textual Time Series: Structuring Narratives for Temporal AI in Healthcare

### v3

@@ Discussion
Appendix C Textual Time Series Forecasting Framework of Encoder Models
Input Preprocessing and Timeseries Construction.
Model Input Representation.
Language Models.
  • bert-base-uncased (BERT; Devlin et al., 2018)
  • roberta-base (Liu et al., 2019)
  • microsoft/deberta-v3-small (He et al., 2021)
  • answerdotai/ModernBERT-base and answerdotai/ModernBERT-large
  • thomas-sounack/BioClinical-ModernBERT-base and thomas-sounack/BioClinical-ModernBERT-large
Prediction Tasks.
1. Event Ordering (Concordance) Task.
  • Input: The serialized history segment followed by:
  • Label: $y=1$ if $\tau_{i}<\tau_{j}$, else $y=0$
2. Time-Window Classification Task.
  • Input: The serialized history segment followed by:
    $\texttt{[SEP]}\;\tau_{j}\texttt{:}f_{j}\;\texttt{[SEP]}$ (3)
Masked Language Modeling Approaches.
1. Masked Event Occurrence Prediction.
  • Target: The model predicts either ”yes” or ”no” at the [MASK] position
2. Masked Temporal Ordering.
  • Input: History followed by: [SEP] $e_{1}$ [MASK] $e_{2}$ [SEP]
Implementation Details.
Training and Evaluation.
Appendix D Survival Modeling with Language Model Embeddings
Table A4: Hyperparameter search ranges for survival models.
Appendix E Computing Resources
All experiments were conducted on a high-performance computing cluster running AlmaLinux. For the largest language models (DSL33 and L33), training required three NVIDIA A100 GPUs (80 GB each; total 240 GB GPU memory), while inference was performed on two A100 GPUs. The encoder models were trained and evaluated on a single NVIDIA A100 GPU (80 GB). After getting the embedding vectors from LLMs, survival models were trained and evaluated on a workstation equipped with an NVIDIA RTX 4090 GPU. System memory usage was 32 GB RAM. Software dependencies, including specific versions of Python libraries and frameworks, are documented in environment_tts_forecasting.yml file of our anonymized repository in the Supplementary Material.
F  LLM prompt to generate clinical textual time-series from PMOA case reports
Prompt You are a physician. Extract the clinical events and the related time stamp from the case report. The admission event has timestamp 0. If the event is not available, we treat the event, e.g. current main clinical diagnosis or treatment with timestamp 0. The events that happened before event with 0 timestamp have negative time, the ones after the event with 0 timestamp have positive time. The timestamp are in hours. The unit will be omitted when output the result. If there is no temporal information of the event, please use your knowledge and events with temporal expression before and after the events to provide an approximation. We want to predict the future events given the events happened in history. For example, here is the case report. An 18-year-old male was admitted to the hospital with a 3-day history of fever and rash. Four weeks ago, he was diagnosed with acne and received the treatment with minocycline, 100 mg daily, for 3 weeks. With increased WBC count, eosinophilia, and systemic involvement, this patient was diagnosed with DRESS syndrome. The fever and rash persisted through admission, and diffuse erythematous or maculopapular eruption with pruritus was present. One day later the patient was discharged.
18 years old | 0
male | 0
admitted to the hospital | 0
fever | -72
rash | -72
acne | -672
minocycline | -672
increased WBC count | 0
eosinophilia | 0
systemic involvement | 0
diffuse erythematous or maculopapular eruption | 0
pruritis | 0
DRESS syndrome | 0
fever persisted | 0
rash persisted | 0
discharged | 24
G  Prompting strategies for forecasting tasks
F.1  Zero-shot
Prompt You are an expert physician. Reply to the prompt with structured predictions in a k-item, bar-separated row. For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response. Reply with the answer only. Do NOT provide any other text. Do NOT explain.
F.2  Zero-shot (with degree of confidence)
Prompt You are an expert physician. Reply to the prompt with structured predictions in a k-item, bar-separated row. Assess whether the event happens in the time window using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen. For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: $1|5|1$ would be the correct response. Reply with the answer only. Do NOT provide any other text. Do NOT explain.
F.3  Few-shot
Prompt You are an expert physician. Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window. Example 1. If there k=3 events (A, B, C) and only B occurs in the requested time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response. Example 2. Suppose the time series is: admitted | 0
history of diabetes | 0
elevated troponin | 5
--- PREDICTION WINDOW ---
cardiac catheterization | 6
admission to icu | 24
transfer to floor | 72
discharge | 144
office visit | 1440
F.4  Few-shot ordinal
Prompt You are an expert physician. Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window, using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen. Example 1. If there k=3 events (A, B, C) and only B occurs in the time window, then: 1 $\mid$ 5 $\mid$ 1 would be the correct response. Example 2. Suppose the time series is: admitted | 0
history of diabetes | 0
elevated troponin | 5
--- PREDICTION WINDOW ---
cardiac catheterization | 6
admission to icu | 24
transfer to floor | 72
discharge | 144
office visit | 1440
Appendix H Open-source models for avoiding potential data leakage
We recognize that data leakage is a critical concern in evaluating model performance, particularly when using open-access text corpora. Below, we outline key factors that mitigate this concern in our study:
1. Temporal annotations—not raw text—drive forecasting tasks. The core contribution of this work lies in the temporal structuring of clinical narratives. While large language models (LLMs) such as DeepSeek and Llama-3.3 may have encountered PubMed Open Access (PMOA) case reports during pretraining, they were never exposed to the extracted timestamps or temporally ordered and future-masked (event, time) tuples used in our evaluations. These derived temporal annotations are essential to all forecasting tasks in our study—including F1-based next-event prediction, temporal ordering, and survival analysis—and constitute novel structured data not present in the original text.
2. Temporal reordering fundamentally alters the learning problem. Even if LLMs were pretrained on PMOA narratives, they would have encountered events in their original, text-based order. In contrast, our framework chronologically reorders events based on extracted timestamps, creating a new temporal structure. This transformation introduces a distinct modeling challenge, as it breaks the sequential narrative flow seen during pretraining. As shown in Section 4.4.1, training on time-ordered versus text-ordered annotations leads to measurable differences in performance, underscoring the impact of this restructuring.
3. External validation reveals limited effects of memorization. Our experimental design separates internal (T2S2 test set) and external validation (sepsis-10 and sepsis-100 cohorts), using entirely held-out case reports. The performance drop observed on external datasets (Tables 1–2) suggests that any potential memorization of PMOA case reports during pretraining yields minimal benefit and reinforces the robustness of our empirical findings.
4. Broader context: standard practices and safeguards. Training data overlap is a general challenge in modern machine learning, especially for LLMs trained on web-scale corpora. Our work adheres to community standards while incorporating additional safeguards—such as explicit temporal annotation, event reordering, and external dataset validation—that reduce the risk of leakage and differentiate our tasks from standard text modeling.
To address concerns regarding potential pretraining overlap, we evaluated multiple models with documented training corpora. Both OLMO-32B-Instruct (Allen Institute for AI) and RedPajama-INCITE-7B-Instruct represent fully open-source models with transparent training data. Notably, RedPajama explicitly excludes PubMed from its training corpus, making it an ideal benchmark for assessing causal leakage. The comparable performance of these models to those potentially exposed to PMOA (see point 5 above) provides strong evidence that our temporal forecasting tasks are meaningfully distinct from standard language modeling on clinical narratives.
5. Empirical evidence from PubMed-excluded models. Our results provide direct empirical evidence against meaningful causal leakage. RedPajama-INCITE-7B-Instruct, which explicitly excludes PubMed from its training corpus, achieves comparable performance to other decoder models (0.352 F1 at 24h, 0.618 concordance) that potentially had access to PMOA during pretraining. This similarity in performance across models with and without PubMed exposure demonstrates that our T2S2 prediction task—with its temporal restructuring, event masking, and partial history inputs—is sufficiently distinct from raw case report text that prior exposure provides no meaningful advantage. Furthermore, OLMO-32B-Instruct shows competitive performance (0.411 F1 at 24h), providing additional validation from a model with fully documented training data.
Appendix I Broader Impact

### v5

@@ H  Sensitivity Analysis of Time-Ordered versus Text-Ordered Training
Appendix J Additional Insights into Forecasting Results
In this section, we provide some additional insights into the forecasting results. Specifically, we summarize two observations that clarify model behavior and apparent anomalies in the results.
Why decoders fare better on survival but not short-horizon forecasting
Our findings align with adjacent literature on representation usage over different time horizons. Encoder models, typically pooled via a [CLS] token, excel on short-horizon tasks because bidirectional attention aggregates local context effectively (Devlin et al. 2019; Bai, Kolter, and Koltun 2018). In contrast, decoder models trained autoregressively pass information forward through final token states, which is advantageous for long-range aggregation required by survival outcomes (Dai et al. 2019; Zhou et al. 2021). In our setting, forecasting aligns with near-term prediction, whereas survival analysis emphasizes longer-range temporal dependencies; this mismatch naturally yields the observed performance split. While not a one-to-one architectural proof, the pattern is consistent with prior evidence on horizon-specific strengths (Wang and Weiss 2025).
Interpreting very low $F_{1}$ scores and external validation drops
Why some $F_{1}$ scores are near zero. The very low F1 scores observed for certain models are not random failures but arise from two interacting factors. First, F1 is highly sensitive to threshold choice when the positive class is rare; small shifts in threshold calibration can yield 0.000 even when the model produces meaningful probabilities. Second, **architectural limitations** amplify this effect: **autoregressive decoders** tend to reproduce narrative rather than temporal order, and **instruction-tuned models** often default to generic or high-probability completions rather than selecting the correct event token—a behavior documented in prior analyses of instruction-following LLMs ((Liu et al. 2024b; Wei et al. 2021)).
These failures do not undermine model comparisons; rather, they highlight systematic architectural differences. **Encoder-based models**, trained with a **bidirectional masking objective** ((Devlin et al. 2019)), are better aligned with time-structured prediction tasks, as their contextual representation naturally integrates information from both past and future tokens. In contrast, **autoregressive decoders**, optimized for left-to-right generation ((Brown et al. 2020)), are disadvantaged for temporal reasoning without explicit sequence restructuring or supervision. The contrast underscores that forecasting from clinical narratives is fundamentally different from standard text generation.
Why external validation is lower. Performance drops on external subsets (e.g., sepsis-10 vs. sepsis-100) reflect shifts in writing style and case mix, not random overfitting. Crucially, the relative ranking of models remains stable across datasets, so comparative conclusions hold even when absolute scores vary.
Appendix K Open-Source Models for Avoiding Potential Data Leakage
We evaluated potential data leakage arising from large language models (LLMs) pretrained on web-scale corpora that may include PubMed Open Access (PMOA) case reports. Our safeguards target two avenues of leakage: (i) content overlap with PMOA text, and (ii) task leakage from using representations too close to raw prose. Below we summarize the controls embedded in our methods and the additional empirical evidence gathered for this concern.
Task Design Minimizes Leakage
Temporal annotations—not raw text—drive all evaluations. We transform narratives into chronological lists of (event, time) tuples and operate on these derived annotations (future-masked where applicable). Models are never scored on lexical similarity to the raw prose.
Temporal reordering changes the learning problem. Events are reordered by extracted timestamps, breaking the original textual sequence seen during pretraining. This creates a distinct modeling challenge: correct temporal ordering and causal dependencies must be inferred from structured tuples rather than reproduced from prose. As shown in the main text, training/evaluating on time-ordered versus text-ordered sequences yields measurably different performance.
Temporal metrics reduce any advantage from memorization. Evaluation emphasizes temporal reasoning—forecasting F1 at fixed horizons, pairwise concordance (c-index), and AULTC—rather than surface-form similarity. Memorizing text order can even hurt time-alignment metrics because clinical narratives often interleave retrospective commentary where textual order $\neq$ chronological order.
Open-Source, PubMed-Free Baselines
Decoder baseline. RedPajama-INCITE-7B-Instruct is pretrained on English CommonCrawl (67%), C4 (15%), GitHub (4.5%), Wikipedia (4.5%), Books (4.5%), arXiv (2.5%), and StackExchange (2%), explicitly excluding PubMed. In our experiments it performs comparably to decoders that may have seen PMOA during pretraining.
Encoder baselines. We include BERT (BooksCorpus + Wikipedia), RoBERTa (CommonCrawl News, OpenWebText, Stories, Wikipedia), and DeBERTa (RoBERTa-style corpora). None of these encoders include PubMed in their pretraining data.
Post-Cutoff External Validation (Additional Evidence)
To probe encoder leakage further, we evaluated BERT, RoBERTa, and DeBERTa on 115 case reports published after their pretraining cutoffs (post-2020). Results mirror submission trends without degradation, which is inconsistent with a memorization hypothesis.
Model 0h 24h 168h
BERT 0.63 0.62 0.51
RoBERTa 0.54 0.61 0.65
DeBERTa 0.58 0.56 0.53
Table A9: Post-cutoff evaluation on 115 case reports (forecasting F1 at 0h/24h/168h).
We also report competitive performance from OLMO-32B-Instruct (fully documented training data) and parity from the PubMed-excluded RedPajama-INCITE-7B-Instruct relative to decoders potentially exposed to PMOA, further indicating that prior PubMed exposure does not meaningfully advantage our temporally structured tasks.
The key takeaways are
  • Using derived, time-ordered (event, time) tuples and temporal metrics directly targets temporal reasoning and reduces susceptibility to content memorization.
  • PubMed-free baselines and post-cutoff tests provide external checks
  • Empirically, models without PubMed exposure achieve comparable performance, arguing against causal leakage into our forecasting, concordance, and AULTC evaluations.
Appendix L Generalizability Beyond Case Reports: Comparison with other Clinical Sources
Scope and Rationale for using Case Reports
We acknowledge that published case reports differ from routine clinical documentation: they emphasize atypical presentations and are written for publication rather than point-of-care use, so they cannot substitute for validation on EHR notes. Our choice is methodological and pragmatic. Case reports are (i) publicly available, (ii) rich in temporally explicit clinical reasoning, and (iii) sufficiently structured to support extraction of textual time series. This makes them a practical, reproducible testbed for assessing whether forecasting models operate on temporally structured narratives without IRB barriers or restricted data access. We therefore use case reports to establish methodological feasibility and transparency, while recognizing that generalization to clinical operations must be demonstrated on EHR notes.
Position Relative to Structured Clinical Scores
Direct comparison to structured risk scores (e.g., SOFA, SAPS) is infeasible because required variables are absent from case reports. Our contribution is complementary to prior work that integrates text with structured data (Kline et al. 2022; Wang and Weiss 2025): we show that free-text narratives, once transformed into timelines, carry prognostic signal usable for forecasting. Whereas Kline et al. (2022) demonstrated that text can augment structured models, our experiments isolate the incremental value of text when converted into time series, independent of structured covariates. In contexts where structured variables are available, these representations can be fused with established scores to assess additive value; here we focus on the text-only signal because the necessary structured inputs are not present in case reports.
Appendix M Broader Societal Impact